OpenStudy (anonymous):

A rectangular parking lot has a length that is 6 feet less than twice the wideth. If the area of the parking lot is 216 square feet what are the parking lots dimensions?

OpenStudy (ddcamp):

Let L be the length, and W be the width. Things we know: L = 2W - 6 Area = L×W = 216

OpenStudy (ddcamp):

To solve, replace L in the second equation with the value of L from the first equation.

OpenStudy (anonymous):

huh

OpenStudy (ddcamp):

(2W-6)×W=216, you can find W. Once you have W, you can find L

OpenStudy (anonymous):

How am I suppose to find W?

OpenStudy (ddcamp):

Solve the equation (2W-6)×W=216 for W

OpenStudy (ddcamp):